OpenStudy (cenaida):

Solve for x by completing the square: 4x2 + 6x - 6 = 0 A. (-3+√33)/4, (-3-√33)/4 B. (3+√33)/4, (3-√33)/4 C. 3, -2 D. 1, 6

OpenStudy (anonymous):

simply tak out 2 common and then use shreedharacharya

OpenStudy (campbell_st):

leave the common factor in as \[(2x)^2 = 4x^2... \]

OpenStudy (campbell_st):

the 1st term in the binomial is 2x since \[(2x)^2 = 4x^2\] so them middle term is \[2 \times 2x + ? = 6x\] find the value of ? then square it... then you'll have \[(2x + ?)^2 - 6 - ? = 0\] then when you have the perfect square you can then look at solving...

OpenStudy (cenaida):

thanks so the answer is B?

OpenStudy (campbell_st):

oops should be \[(2x + ?)^2 - 6 - ?^2 = 0\]

OpenStudy (campbell_st):

I don't know... I just showed you how to complete the square so you can find a solution.

OpenStudy (cenaida):

but can you check if im right?.............

OpenStudy (campbell_st):

oh... and the answers not B

OpenStudy (campbell_st):

but if you wait long enough someone always gives the answer

OpenStudy (anonymous):